# Spurious-Aware Prototype Refinement for Reliable Out-of-Distribution Detection

### v1

@@ 5.3 Results
Table 2: Comparative performance of post-hoc OOD detection methods on SP-OOD benchmarks. Left: AUROC scores (higher is better); Right: FPR@95 scores (lower is better), all in %. Feature-based methods are indicated in blue, output-based methods in red, gradient-based in black, and hybrid methods in green. For each experiment, the top-performing method is shown in bold, and the second-best is underlined.
@@
Energy[13] $62.0_{{\color[rgb]{.5,.5,.5}\pm 2.6}}$ $45.4_{{\color[rgb]{.5,.5,.5}\pm 3.4}}$ $38.4_{{\color[rgb]{.5,.5,.5}\pm 2.1}}$ $79.9_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $80.6_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $61.3$
MLS[44] $62.2_{{\color[rgb]{.5,.5,.5}\pm 2.3}}$ $45.3_{{\color[rgb]{.5,.5,.5}\pm 3.2}}$ $38.4_{{\color[rgb]{.5,.5,.5}\pm 1.4}}$ $80.2_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $81.9_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $61.6$
KLM[44] $51.2_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $41.7_{{\color[rgb]{.5,.5,.5}\pm 2.5}}$ $57.0_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $74.2_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $79.6_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $60.7$
GNorm[50] $79.5_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $38.0_{{\color[rgb]{.5,.5,.5}\pm 1.3}}$ $46.6_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $74.2_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $\underline{85.2}_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $64.7$
ReAct[48] $72.9_{{\color[rgb]{.5,.5,.5}\pm 3.6}}$ $45.6_{{\color[rgb]{.5,.5,.5}\pm 5.3}}$ $41.3_{{\color[rgb]{.5,.5,.5}\pm 3.1}}$ $80.1_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $83.6_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $64.7$
VIM[49] $79.6_{{\color[rgb]{.5,.5,.5}\pm 2.5}}$ $50.4_{{\color[rgb]{.5,.5,.5}\pm 3.1}}$ $60.7_{{\color[rgb]{.5,.5,.5}\pm 1.7}}$ $78.6_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $77.4_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $69.3$
MDS[11] $90.2_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $\underline{57.8}_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $\underline{91.8}_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $62.9_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $58.4_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $72.2$
@@
KNN[46] $\underline{98.6}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $54.5_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $91.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $79.7_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $77.4_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\underline{80.3}$
@@
SPROD ${\mathbf{98.8}}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ ${\mathbf{61.6}}_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ ${\mathbf{97.4}}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ ${\mathbf{82.1}}_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ ${\mathbf{85.3}}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ ${\mathbf{85.0}}$
@@
MSP[7] $87.9_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $98.7_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $97.3_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $83.8_{{\color[rgb]{.5,.5,.5}\pm 6.5}}$ $74.1_{{\color[rgb]{.5,.5,.5}\pm 1.2}}$ $88.4$
Energy[13] $89.2_{{\color[rgb]{.5,.5,.5}\pm 3.2}}$ $98.6_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $95.5_{{\color[rgb]{.5,.5,.5}\pm 3.1}}$ $84.8_{{\color[rgb]{.5,.5,.5}\pm 8.2}}$ $76.3_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $88.9$
MLS[44] $88.1_{{\color[rgb]{.5,.5,.5}\pm 2.0}}$ $98.8_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $96.7_{{\color[rgb]{.5,.5,.5}\pm 2.0}}$ $84.4_{{\color[rgb]{.5,.5,.5}\pm 7.9}}$ $74.6_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $88.5$
KLM[44] $89.1_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $98.7_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $97.1_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $80.5_{{\color[rgb]{.5,.5,.5}\pm 7.6}}$ $76.1_{{\color[rgb]{.5,.5,.5}\pm 1.7}}$ $88.3$
GNorm[50] $84.2_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $98.8_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $97.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $84.2_{{\color[rgb]{.5,.5,.5}\pm 5.8}}$ $\underline{54.7}_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $83.8$
ReAct[48] $86.9_{{\color[rgb]{.5,.5,.5}\pm 7.0}}$ $96.3_{{\color[rgb]{.5,.5,.5}\pm 2.4}}$ $95.5_{{\color[rgb]{.5,.5,.5}\pm 3.2}}$ $83.9_{{\color[rgb]{.5,.5,.5}\pm 8.3}}$ $57.5_{{\color[rgb]{.5,.5,.5}\pm 1.6}}$ $84.0$
VIM[49] $61.4_{{\color[rgb]{.5,.5,.5}\pm 3.5}}$ $96.2_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $69.0_{{\color[rgb]{.5,.5,.5}\pm 1.5}}$ $86.6_{{\color[rgb]{.5,.5,.5}\pm 7.4}}$ $79.5_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $78.5$
MDS[11] $49.2_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $96.0_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $\underline{39.0}_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $93.0_{{\color[rgb]{.5,.5,.5}\pm 3.0}}$ $90.5_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $73.5$
RMDS[45] $91.7_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $99.6_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $95.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $83.4_{{\color[rgb]{.5,.5,.5}\pm 9.3}}$ $88.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $91.6$
KNN[46] $\underline{4.8}_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $\underline{94.4}_{{\color[rgb]{.5,.5,.5}\pm 1.0}}$ $42.5_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $\underline{79.9}_{{\color[rgb]{.5,.5,.5}\pm 10.9}}$ $70.4_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $\underline{58.4}$
SHE[47] $33.2_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $96.4_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $76.5_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $93.9_{{\color[rgb]{.5,.5,.5}\pm 2.9}}$ $\textbf{52.6}_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $70.5$
SPROD $\textbf{4.7}_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $\textbf{93.7}_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $\textbf{19.0}_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $\textbf{70.1}_{{\color[rgb]{.5,.5,.5}\pm 12.0}}$ $58.0_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ 49.1
@@
Zero-shot OOD detection methods have reported results on Waterbirds using CLIP-B/16, but since they operate in a zero-shot setting, spurious correlations in the training set are not meaningful for them. Still, we compare: MCM [29] and CMA [32] achieve 98.36 and 98.62 AUROC using text inputs, while our text-free, vision-only method outperforms both with 99.01. COVER [28] also evaluates this setting but over a different operating range, achieving lower AUROC scores (90.52 vs. 90.31 for MCM), highlighting our method’s superior performance.
The primary results, summarized in Table 2, demonstrate the efficacy of SPROD across various datasets using both AUROC and FPR@95 metrics. SPROD consistently achieves superior performance compared to the 11 post-hoc baselines in all conducted experiments. In contrast, competing methods exhibit more variable performance, excelling in only a subset of the experimental settings. Generally, feature-based OOD detection methods tend to outperform output-based methods; however, this advantage appears to diminish on larger-scale datasets with multiple classes. An interesting observation arises from the Spurious ImageNet results, where SHE achieves the best FPR@95, despite ranking fifth based on AUROC. Overall, the average performance across all datasets reveals a notable margin by which SPROD surpasses the other evaluated baselines: Specifically, SPROD reaches an average of 85.0% AUROC, 4.7% higher (absolute) than the second best, KNN, and 12.8% higher than the third, MDS. Similarly, for FPR@95, SPROD reaches 49.1% error rate on average, 9.3% better than the second best, KNN. All other compared approaches are more than 20% worse. This underscores SPROD’s robustness in detecting SP-OOD samples.
Next, we investigate the impact of backbone fine-tuning and correlation rate on OOD detection performance, using the Waterbirds dataset with ResNet-50 (Figure 4(a)) and ResNet-18 (Figure 4(b)) backbones. A significant observation is that fine-tuning the backbone on this dataset generally degrades OOD detection performance, a finding that contrasts with some common assumptions. This negative effect appears to be more pronounced for feature-based methods. As expected, increasing the spurious correlation rate of ID data from 50% to 90% leads to a general decline in performance across methods; however, this degradation is more noticeable for output-based techniques. Furthermore, the results suggest that employing a lighter, less expressive backbone (ResNet-18 compared to ResNet-50) does not lead to a substantial performance drop in OOD detection. Importantly, across these variations, SPROD demonstrates consistent robustness, particularly maintaining its performance advantage as the spurious correlation rate increases.
@@
To further investigate the impact of the scoring mechanism, as discussed in Section 3.2, we conduct an ablation study using SPROD as the base method. This controlled experiment compares the effectiveness of deriving OOD scores in a discriminative manner $p(y|z)$ versus a generative manner $p(z|y)$. Both baselines utilize the same samples, feature embeddings, and refined prototypes from SPROD. For the discriminative score, we apply a softmax function to the negative distances between a sample’s embedding $z$ and the class prototypes. The default SPROD approach, which uses the negative of the minimum distance to class prototypes, serves as the generative baseline, aligning with a log-likelihood under an exponential family distribution assumption.
Figure 6(a) presents histograms of the scores generated by both approaches, showing that the generative scoring method yields more distinctly separated distributions for ID and OOD samples. Furthermore, Figure 6(b) compares the OOD detection performance of these two scoring variants across our five SP-OOD benchmark settings. The results indicate that applying the softmax function for discriminative scoring substantially degrades performance, particularly in binary classification datasets. Conversely, on datasets with a larger number of classes, such as Animals MetaCoCo and Spurious ImageNet, the performance degradation from using softmax is less pronounced. This observation aligns with the trends in Table 2, where traditional output-based methods tend to perform relatively better in multi-class scenarios compared to binary ones.
@@
6 Discussion, Limitations, and Broader Impact

### v2

@@ 5.3 Results
Table 2: Comparative performance of post-hoc OOD detection methods on SP-OOD benchmarks using a ResNet-50 backbone. Left: AUROC scores (higher is better); Right: FPR@95 scores (lower is better). Feature-based methods are indicated in blue, output-based methods in red, gradient-based in black, and hybrid methods in green. For each experiment, the top-performing method is shown in bold, and the second-best is underlined.
[…]
@@
Energy[13] $62.0_{{\color[rgb]{.5,.5,.5}\pm 2.6}}$ $45.4_{{\color[rgb]{.5,.5,.5}\pm 3.4}}$ $38.4_{{\color[rgb]{.5,.5,.5}\pm 2.1}}$ $79.9_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $80.6_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $61.3$
MLS[44] $62.2_{{\color[rgb]{.5,.5,.5}\pm 2.3}}$ $45.3_{{\color[rgb]{.5,.5,.5}\pm 3.2}}$ $38.4_{{\color[rgb]{.5,.5,.5}\pm 1.4}}$ $80.2_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $81.9_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $61.6$
KLM[44] $51.2_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $41.7_{{\color[rgb]{.5,.5,.5}\pm 2.5}}$ $57.0_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $74.2_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $79.6_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $60.7$
GEN[45] $62.3_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $46.0_{{\color[rgb]{.5,.5,.5}\pm 1.4}}$ $38.5_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $2_{{\color[rgb]{.5,.5,.5}\pm 0.0}} 80.8±0.$$8_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $61.6$
GNorm[61] $79.5_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $38.0_{{\color[rgb]{.5,.5,.5}\pm 1.3}}$ $46.6_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $74.2_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $85.2_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $64.7$
ReAct[58] $72.9_{{\color[rgb]{.5,.5,.5}\pm 3.6}}$ $45.6_{{\color[rgb]{.5,.5,.5}\pm 5.3}}$ $41.3_{{\color[rgb]{.5,.5,.5}\pm 3.1}}$ $80.1_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $83.6_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $64.7$
VIM[59] $79.6_{{\color[rgb]{.5,.5,.5}\pm 2.5}}$ $50.4_{{\color[rgb]{.5,.5,.5}\pm 3.1}}$ $60.7_{{\color[rgb]{.5,.5,.5}\pm 1.7}}$ $8.6_{{\color[rgb]{.5,.5,.5}\pm 0.6}} 77$$7.4_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $69.3$
ASH[60] $78.5_{{\color[rgb]{.5,.5,.5}\pm 3.2}}$ $47.3_{{\color[rgb]{.5,.5,.5}\pm 2.8}}$ $39.6_{{\color[rgb]{.5,.5,.5}\pm 1.7}}$ $78.0_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $\textbf{86.6}_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $66.0$
MDS[11] $90.2_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $57.8_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $91.8_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $62.9_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $58.4_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $72.2$
@@
KNN[47] $\underline{98.6}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $54.5_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $91.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $9.7_{{\color[rgb]{.5,.5,.5}\pm 0.6}} 77$$7.4_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\underline{80.3}$
@@
NECO[51] $53.5_{{\color[rgb]{.5,.5,.5}\pm 1.6}}$ $39.5_{{\color[rgb]{.5,.5,.5}\pm 3.2}}$ $35.1_{{\color[rgb]{.5,.5,.5}\pm 1.5}}$ $80.2_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $67.2_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $55.1$
NNGuide[49] $70.6_{{\color[rgb]{.5,.5,.5}\pm 2.9}}$ $49.8_{{\color[rgb]{.5,.5,.5}\pm 4.2}}$ $43.6_{{\color[rgb]{.5,.5,.5}\pm 2.1}}$ $79.4_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $85.1_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $65.7$
Relation[50] $80.7_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $\underline{60.4}_{{\color[rgb]{.5,.5,.5}\pm 2.5}}$ $\underline{96.0}_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $74.5_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $81.8_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $78.7$
SCALE[52] $89.0_{{\color[rgb]{.5,.5,.5}\pm 2.9}}$ $44.9_{{\color[rgb]{.5,.5,.5}\pm 3.2}}$ $54.4_{{\color[rgb]{.5,.5,.5}\pm 2.1}}$ $78.4_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $\underline{86.2}_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $70.6$
fDBD[53] $71.1_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $51.3_{{\color[rgb]{.5,.5,.5}\pm 1.3}}$ $47.4_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $79.9_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $84.2_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $66.8$
NCI[54] $84.0_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $46.4_{{\color[rgb]{.5,.5,.5}\pm 2.4}}$ $54.8_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $78.5_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $84.9_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $69.7$
SPROD $\textbf{98.8}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{61.6}_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $\textbf{97.4}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $\textbf{82.4}_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $85.3_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ 85.1
[…]
@@
MSP[7] $87.9_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $98.7_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $97.3_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $83.8_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $74.1_{{\color[rgb]{.5,.5,.5}\pm 1.2}}$ $88.4$
Energy[13] $89.2_{{\color[rgb]{.5,.5,.5}\pm 3.2}}$ $98.6_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $95.5_{{\color[rgb]{.5,.5,.5}\pm 3.1}}$ $84.8_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $76.3_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $88.9$
MLS[44] $88.1_{{\color[rgb]{.5,.5,.5}\pm 2.0}}$ $98.8_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $96.7_{{\color[rgb]{.5,.5,.5}\pm 2.0}}$ $84.4_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $74.6_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $88.5$
KLM[44] $89.1_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $98.7_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $97.1_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $80.5_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $76.1_{{\color[rgb]{.5,.5,.5}\pm 1.7}}$ $88.3$
GEN[45] $87.9_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $98.7_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $97.3_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $84.8_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $76.3_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $89.0$
GNorm[61] $84.2_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $98.8_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $97.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $84.2_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $54.7_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $83.8$
ReAct[58] $86.9_{{\color[rgb]{.5,.5,.5}\pm 7.0}}$ $96.3_{{\color[rgb]{.5,.5,.5}\pm 2.4}}$ $95.5_{{\color[rgb]{.5,.5,.5}\pm 3.2}}$ $83.9_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $57.5_{{\color[rgb]{.5,.5,.5}\pm 1.6}}$ $84.0$
VIM[59] $61.4_{{\color[rgb]{.5,.5,.5}\pm 3.5}}$ $96.2_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $69.0_{{\color[rgb]{.5,.5,.5}\pm 1.5}}$ $86.6_{{\color[rgb]{.5,.5,.5}\pm 0.7}}$ $79.5_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $78.5$
ASH[60] $85.2_{{\color[rgb]{.5,.5,.5}\pm 7.0}}$ $96.9_{{\color[rgb]{.5,.5,.5}\pm 1.4}}$ $96.1_{{\color[rgb]{.5,.5,.5}\pm 1.5}}$ $87.9_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $52.9_{{\color[rgb]{.5,.5,.5}\pm 3.1}}$ $83.8$
MDS[11] $49.2_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $96.0_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $_{{\color[rgb]{.5,.5,.5}\pm 0.3}} 93.0±0.39$$_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $90.5_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $73.5$
RMDS[46] $1.7_{{\color[rgb]{.5,.5,.5}\pm 0.2}} 99$$9.6_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $95.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $3.4_{{\color[rgb]{.5,.5,.5}\pm 0.9}} 88$$8.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $91.6$
KNN[47] $\underline{4.8}_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $\underline{94.4}_{{\color[rgb]{.5,.5,.5}\pm 1.0}}$ $42.5_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $79.9_{{\color[rgb]{.5,.5,.5}\pm 1.1}}$ $70.4_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $\underline{58.4}$
SHE[48] $33.2_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $96.4_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $76.5_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $93.9_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $\underline{52.6}_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $70.5$
NECO[51] $90.5_{{\color[rgb]{.5,.5,.5}\pm 2.2}}$ $98.8_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $96.7_{{\color[rgb]{.5,.5,.5}\pm 1.9}}$ $\underline{78.2}_{{\color[rgb]{.5,.5,.5}\pm 0.0}}$ $89.9_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $90.8$
NNGuide[49] $77.7_{{\color[rgb]{.5,.5,.5}\pm 6.0}}$ $97.6_{{\color[rgb]{.5,.5,.5}\pm 1.2}}$ $91.6_{{\color[rgb]{.5,.5,.5}\pm 3.6}}$ $86.3_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $\textbf{52.2}_{{\color[rgb]{.5,.5,.5}\pm 1.4}}$ $81.1$
Relation[50] $73.8_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $95.4_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $\underline{24.2}_{{\color[rgb]{.5,.5,.5}\pm 1.7}}$ $84.6_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $78.0_{{\color[rgb]{.5,.5,.5}\pm 1.1}}$ $71.2$
SCALE[52] $61.0_{{\color[rgb]{.5,.5,.5}\pm 22.5}}$ $98.7_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $94.6_{{\color[rgb]{.5,.5,.5}\pm 3.1}}$ $87.7_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $53.0_{{\color[rgb]{.5,.5,.5}\pm 1.5}}$ $79.0$
fDBD[53] $85.5_{{\color[rgb]{.5,.5,.5}\pm 0.8}}$ $98.6_{{\color[rgb]{.5,.5,.5}\pm 0.5}}$ $96.1_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $85.4_{{\color[rgb]{.5,.5,.5}\pm 0.2}}$ $70.8_{{\color[rgb]{.5,.5,.5}\pm 1.0}}$ $87.3$
NCI[54] $41.1_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $99.4_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $92.2_{{\color[rgb]{.5,.5,.5}\pm 0.6}}$ $85.8_{{\color[rgb]{.5,.5,.5}\pm 0.3}}$ $63.8_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $76.5$
SPROD $\textbf{4.7}_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ $\textbf{93.7}_{{\color[rgb]{.5,.5,.5}\pm 0.9}}$ $\textbf{19.0}_{{\color[rgb]{.5,.5,.5}\pm 0.4}}$ $\textbf{69.5}_{{\color[rgb]{.5,.5,.5}\pm 1.2}}$ $58.0_{{\color[rgb]{.5,.5,.5}\pm 0.1}}$ 49.0
@@
The primary results, summarized in Table 2, demonstrate the efficacy of SPROD across various datasets using a ResNet-50 backbone, evaluated with both AUROC and FPR@95 metrics. Additional results on other backbone architectures are provided in Appendix G. SPROD consistently achieves superior performance compared to the 19 post-hoc baselines in all conducted experiments. In contrast, competing methods exhibit more variable performance, excelling in only a subset of the experimental settings. Generally, feature-based OOD detection methods tend to outperform output-based methods; however, this advantage appears to diminish on larger-scale datasets with multiple classes. Overall, the average performance across all datasets reveals a notable margin by which SPROD surpasses the other evaluated baselines: Specifically, SPROD reaches an average of 85.1% AUROC, 4.8% higher (absolute) than the second best, KNN. For FPR@95, SPROD reaches 49.1% error rate on average, 9.3% better than the second best, KNN. All other compared approaches are more than 20% worse.
Next, we investigate the impact of backbone fine-tuning and correlation rate on OOD detection performance, using the Waterbirds dataset with ResNet-50 (Figure 4(a)) and ResNet-18 (Figure 4(b)) backbones. Fine-tuning the backbone on this dataset generally degrades OOD detection performance, a finding that contrasts with some common assumptions. This negative effect appears to be more pronounced for feature-based methods. As expected, increasing the spurious correlation rate of ID data from 50% to 90% leads to a general decline in performance across methods; however, this degradation is more noticeable for output-based techniques. Furthermore, the results suggest that employing a lighter, less expressive backbone (ResNet-18 compared to ResNet-50) does not lead to a substantial performance drop in OOD detection. Across these variations, SPROD demonstrates consistent robustness and maintains its performance as the spurious correlation rate increases.
@@
To investigate the impact of the scoring mechanism, as discussed in Section 3.2, we conduct an ablation study using SPROD as the base method. This controlled experiment compares the effectiveness of deriving OOD scores in a discriminative manner $p(y|z)$ versus a generative manner $p(z|y)$. Both baselines utilize the same samples, feature embeddings, and refined prototypes from SPROD. For the discriminative score, we apply a softmax function to the negative distances between a sample’s embedding $z$ and the class prototypes. The default SPROD approach, which uses the negative of the minimum distance to class prototypes, serves as the generative baseline (aligning with a log-likelihood under an exponential family distribution assumption).
Figure 6(a) presents histograms of the scores generated by both approaches, showing that the generative scoring method yields more distinctly separated distributions for ID and OOD samples. Furthermore, Figure 6(b) compares the OOD detection performance of these two scoring variants across our five SP-OOD benchmark settings. The results indicate that applying the softmax function for discriminative scoring substantially degrades performance, particularly in WaterBirds, CelebA and UrbanCars datasets. Conversely, on Animals MetaCoCo and Spurious ImageNet, the performance degradation from using softmax is less pronounced. This observation aligns with the trends in Table 2, where traditional output-based methods tend to perform relatively better in these datasets.
[…]
Method CIFAR-10 Near CIFAR-10 Far CIFAR-100 Near CIFAR-100 Far ImageNet Near ImageNet Far Avg (%)
RMDS [46] 89.80 92.20 80.15 82.92 76.99 86.38 84.74
MLS [44] 87.52 91.10 81.05 79.67 76.46 89.57 84.23
VIM [59] 88.68 93.48 74.98 81.70 72.08 92.68 83.27
KNN [47] 90.64 92.96 80.18 82.40 71.10 90.18 84.58
ASH [60] 75.12 78.49 78.20 80.58 78.17 95.74 81.05
SPROD 89.04 91.78 81.80 79.93 75.06 95.29 85.15
To further assess the generality of SPROD beyond spurious correlation settings, we evaluate it on conventional OOD benchmarks using the standardized OpenOOD protocol [37]. We compare SPROD with the top-performing post-hoc methods reported in the OpenOOD study. Table 3 shows that SPROD attains performance comparable to or exceeding these methods and achieves the highest average score across all settings, confirming that SPROD performs reliably on standard OOD benchmarks and is not limited to spurious correlation scenarios.
@@
Zero-shot OOD detection methods have reported results on Waterbirds using CLIP-B/16, but since they operate in a zero-shot setting, spurious correlations in the training set are not meaningful for them. Still, we compare: MCM [29] and CMA [32] achieve 98.36 and 98.62 AUROC using text inputs, while our text-free, vision-only method outperforms both with 99.01. COVER [28] also evaluates this setting but over a different operating range, achieving lower AUROC scores (90.52 vs. 90.31 for MCM), highlighting our method’s superior performance. Additional comparisons with existing SP-OOD methods (specifically CLIP-based approaches) are provided in Appendix C.
6 Discussion